Given f(x) and g(x), find the following. 3) f(x) = x2 + 2x and g(x)= x + 9. (fg)(x)
Answer #1

Soni- Given! - X = x2 + 2x and ; (32) = 279 Then Flagg () 3 = ? = f(2) 9 (0) = (3x2+2x)(x+9) 13 + 2x² + 9x² + 18x 4g)(x) x3 +